OpenStudy (anonymous):

Thirteen cards are dealt from a well shuffled standard deck. Find probability of getting all hearts.

OpenStudy (misty1212):

HI!! close enuf to zero that it might as well be zero

OpenStudy (misty1212):

\[\frac{13\times 12\times 11\times ...\times 1}{52\times 51\times 50\times ...\times 40}\]

OpenStudy (perl):

note that you are drawing the cards without replacement. so the first heart has probability of 13/52 . the second heart , now you only have 12 hearts and 51 cards to choose from , so probability is 12/51 for drawing a heart given that you already drew a heart. and you continue this process, the multiplication rule states we multiply these probabilities